\begin{thebibliography}{25}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Bengio et~al.(2003)Bengio, Ducharme, Vincent, and
  Jauvin}]{bengio2003neural}
Yoshua Bengio, R{\'e}jean Ducharme, Pascal Vincent, and Christian Jauvin. 2003.
\newblock A neural probabilistic language model.
\newblock \emph{Journal of machine learning research}, 3(Feb):1137--1155.

\bibitem[{Conneau et~al.(2017)Conneau, Lample, Ranzato, Denoyer, and
  J{\'e}gou}]{conneau2017word}
Alexis Conneau, Guillaume Lample, Marc'Aurelio Ranzato, Ludovic Denoyer, and
  Herv{\'e} J{\'e}gou. 2017.
\newblock Word translation without parallel data.
\newblock \emph{arXiv preprint arXiv:1710.04087}.

\bibitem[{{Deutsches Textarchiv}(2017)}]{dta2017}
{Deutsches Textarchiv}.
\newblock {Grundlage fuer ein Referenzkorpus der neuhochdeutschen Sprache.
  Herausgegeben von der Berlin-Brandenburgischen Akademie der Wissenschaften}
  [online]. 2017.

\bibitem[{Dubossarsky et~al.(2019)Dubossarsky, Hengchen, Tahmasebi, and
  Schlechtweg}]{DBLP:journals/corr/abs-1906-01688}
Haim Dubossarsky, Simon Hengchen, Nina Tahmasebi, and Dominik Schlechtweg.
  2019.
\newblock Time-out: Temporal referencing for robust modeling of lexical
  semantic change.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 457--470, Florence, Italy. Association
  for Computational Linguistics.

\bibitem[{Dubossarsky et~al.(2017)Dubossarsky, Weinshall, and
  Grossman}]{dubossarsky2017}
Haim Dubossarsky, Daphna Weinshall, and Eitan Grossman. 2017.
\newblock Outta control: Laws of semantic change and inherent biases in word
  representation models.
\newblock In \emph{{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}}, pages 1147--1156, Copenhagen, Denmark.

\bibitem[{Faruqui et~al.(2015)Faruqui, Tsvetkov, Yogatama, Dyer, and
  Smith}]{DBLP:journals/corr/FaruquiTYDS15}
Manaal Faruqui, Yulia Tsvetkov, Dani Yogatama, Chris Dyer, and Noah~A. Smith.
  2015.
\newblock Sparse overcomplete word vector representations.
\newblock \emph{CoRR}, abs/1506.02004.

\bibitem[{Frermann and Lapata(2016)}]{frermann2016bayesian}
Lea Frermann and Mirella Lapata. 2016.
\newblock A bayesian model of diachronic meaning change.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  4:31--45.

\bibitem[{Hamilton et~al.(2016)Hamilton, Leskovec, and
  Jurafsky}]{hamilton-etal-2016-diachronic}
William~L. Hamilton, Jure Leskovec, and Dan Jurafsky. 2016.
\newblock Diachronic word embeddings reveal statistical laws of semantic
  change.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1489--1501,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Joulin et~al.(2016)Joulin, Grave, Bojanowski, Douze, J{\'e}gou, and
  Mikolov}]{joulin2016fasttext}
Armand Joulin, Edouard Grave, Piotr Bojanowski, Matthijs Douze, H{\'e}rve
  J{\'e}gou, and Tomas Mikolov. 2016.
\newblock Fasttext.zip: Compressing text classification models.
\newblock \emph{arXiv preprint arXiv:1612.03651}.

\bibitem[{Kim et~al.(2014)Kim, Chiu, Hanaki, Hegde, and
  Petrov}]{kim-etal-2014-temporal}
Yoon Kim, Yi-I Chiu, Kentaro Hanaki, Darshan Hegde, and Slav Petrov. 2014.
\newblock Temporal analysis of language through neural language models.
\newblock In \emph{Proceedings of the {ACL} 2014 Workshop on Language
  Technologies and Computational Social Science}, pages 61--65, Baltimore, MD,
  USA. Association for Computational Linguistics.

\bibitem[{Kulkarni et~al.(2014)Kulkarni, Al{-}Rfou, Perozzi, and
  Skiena}]{DBLP:journals/corr/KulkarniAPS14a}
Vivek Kulkarni, Rami Al{-}Rfou, Bryan Perozzi, and Steven Skiena. 2014.
\newblock Statistically significant detection of linguistic change.
\newblock \emph{CoRR}, abs/1411.3315.

\bibitem[{Lau et~al.(2012)Lau, Cook, McCarthy, Newman, and
  Baldwin}]{lau2012word}
Jey~Han Lau, Paul Cook, Diana McCarthy, David Newman, and Timothy Baldwin.
  2012.
\newblock Word sense induction for novel sense detection.
\newblock In \emph{Proceedings of the 13th Conference of the European Chapter
  of the Association for Computational Linguistics}, pages 591--601.
  Association for Computational Linguistics.

\bibitem[{Lin(1991)}]{lin1991divergence}
Jianhua Lin. 1991.
\newblock Divergence measures based on the shannon entropy.
\newblock \emph{IEEE Transactions on Information theory}, 37(1):145--151.

\bibitem[{Mikolov et~al.(2013)Mikolov, Chen, Corrado, and
  Dean}]{mikolov2013efficient}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013.
\newblock Efficient estimation of word representations in vector space.
\newblock \emph{arXiv preprint arXiv:1301.3781}.

\bibitem[{Mitra et~al.(2015)Mitra, Mitra, Maity, Riedl, Biemann, Goyal, and
  Mukherjee}]{mitra2015automatic}
Sunny Mitra, Ritwik Mitra, Suman~Kalyan Maity, Martin Riedl, Chris Biemann,
  Pawan Goyal, and Animesh Mukherjee. 2015.
\newblock An automatic approach to identify word sense changes in text media
  across timescales.
\newblock \emph{Natural Language Engineering}, 21(5):773--798.

\bibitem[{Pennington et~al.(2014)Pennington, Socher, and
  Manning}]{pennington2014glove}
Jeffrey Pennington, Richard Socher, and Christopher Manning. 2014.
\newblock Glove: Global vectors for word representation.
\newblock In \emph{Proceedings of the 2014 conference on empirical methods in
  natural language processing (EMNLP)}, pages 1532--1543.

\bibitem[{Ramos et~al.(2003)}]{ramos2003using}
Juan Ramos et~al. 2003.
\newblock Using tf-idf to determine word relevance in document queries.
\newblock In \emph{Proceedings of the first instructional conference on machine
  learning}, volume 242, pages 133--142. Piscataway, NJ.

\bibitem[{Rosenfeld and Erk(2018)}]{rosenfeld-erk-2018-deep}
Alex Rosenfeld and Katrin Erk. 2018.
\newblock Deep neural models of semantic shift.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 474--484, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Rudolph and Blei(2017)}]{rudolph2017dynamic}
Maja Rudolph and David Blei. 2017.
\newblock Dynamic bernoulli embeddings for language evolution.
\newblock \emph{arXiv preprint arXiv:1703.08052}.

\bibitem[{Schlechtweg et~al.(2019)Schlechtweg, H{\"a}tty, Del~Tredici, and
  Schulte~im Walde}]{DBLP:journals/corr/abs-1906-02979}
Dominik Schlechtweg, Anna H{\"a}tty, Marco Del~Tredici, and Sabine Schulte~im
  Walde. 2019.
\newblock A wind of change: Detecting and evaluating lexical semantic change
  across times and domains.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 732--746, Florence, Italy. Association
  for Computational Linguistics.

\bibitem[{Schlechtweg et~al.(2018)Schlechtweg, Schulte~im Walde, and
  Eckmann}]{schlechtweg2018diachronic}
Dominik Schlechtweg, Sabine Schulte~im Walde, and Stefanie Eckmann. 2018.
\newblock Diachronic usage relatedness ({DUR}el): A framework for the
  annotation of lexical semantic change.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pages 169--174, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{{Tahmasebi} et~al.(2018){Tahmasebi}, {Borin}, and
  {Jatowt}}]{2018arXiv181106278T}
Nina {Tahmasebi}, Lars {Borin}, and Adam {Jatowt}. 2018.
\newblock {Survey of Computational Approaches to Diachronic Conceptual Change}.
\newblock \emph{arXiv e-prints}.

\bibitem[{Traugott and Dasher(2001)}]{traugott2001regularity}
Elizabeth~Closs Traugott and Richard~B Dasher. 2001.
\newblock \emph{Regularity in semantic change}, volume~97.
\newblock Cambridge University Press.

\bibitem[{Wang and McCallum(2006)}]{wang2006topics}
Xuerui Wang and Andrew McCallum. 2006.
\newblock Topics over time: a non-markov continuous-time model of topical
  trends.
\newblock In \emph{Proceedings of the 12th ACM SIGKDD international conference
  on Knowledge discovery and data mining}, pages 424--433. ACM.

\bibitem[{Yehezkel~Lubin et~al.(2019)Yehezkel~Lubin, Goldberger, and
  Goldberg}]{yehezkel-lubin-etal-2019-aligning}
Noa Yehezkel~Lubin, Jacob Goldberger, and Yoav Goldberg. 2019.
\newblock Aligning vector-spaces with noisy supervised lexicon.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 460--465, Minneapolis,
  Minnesota. Association for Computational Linguistics.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperindex,breaklinks]{hyperref}
\usepackage{breakurl}
\usepackage[]{acl2019}
\usepackage{times}
\usepackage{latexsym}

\aclfinalcopy 

\title{Shared Task: Lexical Semantic Change Detection in German}

\author{Adnan Ahmad, Kiflom Desta, Fabian Lang, Dominik Schlechtweg \\
  Institute for Natural Language Processing, University of Stuttgart\\
  {\tt adnan.sust@outlook.com}, {\tt desta7883@gmail.com},\\ {\tt f.lang.1993@web.de}, {\tt schlecdk@ims.uni-stuttgart.de}  }

\date{}

\begin{document}
\maketitle
\begin{abstract}
Recent NLP architectures have illustrated in various ways how semantic change can be captured across time and domains. However, in terms of evaluation there is a lack of benchmarks to compare the performance of these systems against each other. We present the results of the first shared task on unsupervised lexical semantic change detection (LSCD) in German based on the evaluation framework proposed by \citet{DBLP:journals/corr/abs-1906-02979}.
\end{abstract}

\section{Introduction}
Natural languages evolve and words have always been subject to semantic change over time \cite{traugott2001regularity}. With the rise of large digitized text resources recent NLP technologies have made it possible to capture such change with vector space models \citep*{pennington2014glove,rudolph2017dynamic,bengio2003neural,rosenfeld-erk-2018-deep}, topic models \citep*{wang2006topics,lau2012word,frermann2016bayesian}, and sense clustering models \citep{mitra2015automatic}. However, many approaches for detecting LSC differ profoundly from each other and therefore drawing comparisons between them can be challenging \citep{2018arXiv181106278T}. Not only do architectures for detecting LSC vary, their performance is also often evaluated without access to evaluation data or too sparse data sets. In cases where evaluation data is available, oftentimes LSCD systems are not evaluated on the same data set which hinders the research community to draw comparisons. 

For this reason we report the results of the first shared task on unsupervised lexical semantic change detection in German\footnote{\url{https://codalab.lri.fr/competitions/560}} that is based on an annotated data set to guarantee objective reasoning throughout different approaches. The task was organized as part of the seminar 'Lexical Semantic Change Detection' at the IMS Stuttgart in the summer term of 2019.\footnote{\url{https://www.f05.uni-stuttgart.de/informatik/dokumente/Seminare/Seminare-SS_2019/04_LexicalSemantics.pdf}}

\section{Task}
The goal of the shared task was to create an architecture to detect semantic change and to rank words according to their degree of change between two different time periods. Given two corpora C\textsubscript{a} and C\textsubscript{b}, the target words had to be ranked according to their degree of lexical semantic change between C\textsubscript{a} and C\textsubscript{b} as annotated by human judges. A competition was set up on Codalab and teams mostly consisting of 2 people were formed to take part in the task. There was one group consisting of 3 team members and two individuals who entered the task on their own. In total there were 12 LSCD systems participating in the shared task.

The shared task was divided into three phases, i.e., development, testing and analysis phase. In the development phase each team implemented a first version of their model based on a trial data set and submitted it subsequently. In the testing phase the testing data was made public and participants applied their models to the test data with a restriction of possible result uploads to 30. The leaderboard was public at all times. Eventually, the analysis phase was entered and the models of the testing phase were evaluated in terms of the predictions they made and parameters could be tuned further. The models and results will be discussed in detail in sections 7 and 8.

\section{Corpora}
The task, as framed above, requires to detect the semantic change between two corpora. The two corpora used in the shared task correspond to the diachronic corpus pair from \citet{DBLP:journals/corr/abs-1906-02979}: DTA18 and DTA19.\footnote{\url{https://www.ims.uni-stuttgart.de/forschung/ressourcen/korpora/wocc/}} They consist of subparts of DTA corpus \citep{dta2017} which is a freely available lemmatized, POS-tagged and spelling-normalized diachronic corpus of German containing texts from the 16th to the 20th century. DTA18 contains 26 million sentences published between 1750-1799 and DTA19 40 million between 1850-1899. The corpus version used in the task has the following format: "year [tab] lemma1 lemma2 lemma3 ...".

\section{Evaluation}
The Diachronic Usage Relatedness (DURel) gold standard data set includes 22 target words and their varying degrees of semantic change \cite{schlechtweg2018diachronic}. For each of these target words a random sample of use pairs from the DTA corpus was retrieved and annotated. The annotators were required to rate the pairs according to their semantic relatedness on a scale from 1 to 4 (unrelated - identical meanings) for two time periods. The average Spearman's $\rho$ between the five annotators was 0.66 for 1,320 use paris. The resulting word ranking of the DURel data set is determined by the mean usage relatedness across two time periods and is used as the benchmark to compare the models’ performances in the shared task. 

\subsection{Metric}
The output of a system with the target words in the predicted order is compared to the gold ranking of the DURel data set. As the metric to assess how well the model's output fits the gold ranking Spearman's $\rho$ was used. The higher Spearman's rank-order correlation the better the system's performance.

\subsection{Baselines}
Models were compared to two baselines for the shared task:
\begin{enumerate}
    \item log-transformed normalized frequency difference (FD)
    \item count vectors with column intersection and cosine distance (CNT + CI + CD)
\end{enumerate}
The window size for CNT + CI + CD was 10. Find more information on these models in \citet{DBLP:journals/corr/abs-1906-02979}.

\begin{table*}[]
\begin{tabular}{|l|lll|l|l|}
\hline
\textbf{Team} & \multicolumn{1}{l|}{\textbf{Space}} & \multicolumn{1}{l|}{\textbf{Align}} & \textbf{Measure} & \textbf{Spearman} & \textbf{Comment} \\ \hline
sorensbn & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{OP} & CD & .854 & Noise-aware alignment \\ \hline
tidoe & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{OP} & CD & .811 & Binarized matrices \\ \hline
in vain & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{VI} & CD & .802 &  \\ \hline
Evilly & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{OP} & CD & .730 & Frequency-driven OP alignment \\ \hline
DAF & \multicolumn{1}{l|}{fastText} & \multicolumn{1}{l|}{OP} & CD & .570 &  \\ \hline
SnakesOnAPlane & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{OP} & CD / JSD & .565 / .561 & Measure comparison \\ \hline
TeamKulkarni15 & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{OP} & CD & .540 & Local alignment with KNN \\ \hline
Bashmaistori & \multicolumn{1}{l|}{PPMI} & \multicolumn{1}{l|}{WI} & CD & .511 &  \\ \hline
Baseline 2 & \multicolumn{1}{l|}{CNT} & \multicolumn{1}{l|}{CI} & CD & .486 & \\ \hline
giki & \multicolumn{1}{l|}{PPMI} & \multicolumn{1}{l|}{CI} & CD & .432 &  \\ \hline
Edu-Phil & \multicolumn{1}{l|}{fastText} & \multicolumn{1}{l|}{OP} & CD & .381 &  \\ \hline
orangefoxes & \multicolumn{1}{l|}{SGNS} & \multicolumn{1}{l|}{-} & CD & .121 & DiffTime \\ \hline
Loud Whisper & \multicolumn{1}{l|}{-} & \multicolumn{1}{l|}{-} & - & .092 & Graph-based approach \\ \hline
Baseline 1 & \multicolumn{1}{l|}{-} & \multicolumn{1}{l|}{-} & FD & .019 &  \\ \hline
\end{tabular}
\caption{Shared task - Overview of participating systems. Legend: Space = Semantic space; Align = Alignment method; Measure = Distance Measure for LSC detection.
\textbf{Note:} The best result either obtained in testing or analysis phase is reported.}
\label{table:tab1}
\end{table*}

\section{Participating Systems}
Participants mostly rely on the models compared in \citet{DBLP:journals/corr/abs-1906-02979} and apply modifications to improve them.\footnote{Find implementations at \url{https://github.com/Garrafao/LSCDetection}.} In particular, most teams make use of skip-gram with negative sampling (SGNS) based on \citet{mikolov2013efficient} to learn the semantic spaces of the two time periods and orthogonal procrustes (OP) to align these vector spaces, similar to the approach by \citet{hamilton-etal-2016-diachronic}. Different meaning representations such as sense clusters are used as well. As measure to detect the degree of LSC all teams except one choose cosine distance (CD). This team uses Jensen-Shannon distance (JSD) instead, which computes the distance between probability distributions \cite{lin1991divergence}.

The models of each team will be briefly introduced in this section. \paragraph{sorensbn} Team sorensbn makes use of SGNS + OP + CD to detect LSC. They use similar hyperparameters as in \citet{DBLP:journals/corr/abs-1906-02979} to tune the SGNS model. They use an open-sourced noise-aware implementation to improve the OP alignment \cite{yehezkel-lubin-etal-2019-aligning}. \paragraph{tidoe} Team tidoe builds on SGNS + OP + CD, but they add a transformation step to receive binarized representations of matrices \cite{DBLP:journals/corr/FaruquiTYDS15}. This step is taken to counter the bias that can occur in vector-space models based on frequencies \citep{dubossarsky2017}. \paragraph{in vain} The team applies a model based on SGNS with vector initialization alignment and cosine distance (SGNS + VI + CD). Vector initialization is an alignment strategy where the vector space learning model for $t_2$ is initialized with the vectors from $t_1$ \cite{kim-etal-2014-temporal}. Since SGNS + VI + OP does not perform as well as other models in \citet{DBLP:journals/corr/abs-1906-02979}, they alter the vector initialization process by initializing on the complete model instead of only the word matrix of $t_1$ to receive improved results. \paragraph{Evilly} In line with previous approaches, team Evilly builds upon SGNS + OP + CD. They alter the OP step by using only high-frequency words for alignment. \paragraph{DAF} Team DAF uses an architecture based on learning vectors with fastText, alignment with unsupervised and supervised variations of OP, and CD, using the MUSE package\footnote{Find package at: \url{https://github.com/facebookresearch/MUSE}} \citep*{conneau2017word,joulin2016fasttext}. For the supervised alignment stop words are used. The underlying assumption is that stop words serve as functional units of language and their usage should be consistent over time. \paragraph{SnakesOnAPlane}
The team learns vector spaces with count vectors, positive pointwise mutual information (PPMI), SGNS and uses column intersection (CI) and OP as alignment techniques where applicable. Then they compare two distance measures (CD and JSD) for the different models CNT + CI, PPMI + CI and SGNS + OP to identify which measure performs better for these models. They also experiment with different ways to remove negative values from SGNS vectors, which is needed for JSD. \paragraph{TeamKulkarni15} TeamKulkarni15 uses SGNS + OP + CD with the modification of local alignment with k nearest neighbors, since other models often use global alignment that can be prone to noise \cite{DBLP:journals/corr/KulkarniAPS14a}. \paragraph{Bashmaistori} They use word injection (WI) alignment on PPMI vectors with CD. This approach avoids the complex alignment procedure for embeddings and is applicable to embeddings and count-based methods. They compare two implementations of word injection \citep*{DBLP:journals/corr/abs-1906-01688,DBLP:journals/corr/abs-1906-02979} as these showed different results on different data sets.  \paragraph{giki} Team giki uses PPMI + CI + CD to detect LSC. They state that a word sense is determined by its context, but relevant context words can also be found outside a predefined window. Therefore, they use tf-idf to select relevant context \cite{ramos2003using}. \paragraph{Edu-Phil} Similar to team DAF they also use fastText + OP + CD. Their hypothesis is that fastText may increase the performance for less frequent words in the corpus since generating word embeddings in fasttext is based on character n-grams. \paragraph{orangefoxes} They use the model by \citet{rosenfeld-erk-2018-deep} which is based on SGNS, but avoids alignment by treating time as a vector that may be combined with word vectors to get time-specific word vectors. \paragraph{Loud Whisper} Loud Whisper base their approach on \citet{mitra2015automatic} which is a graph-based sense clustering model. They process the data set to receive bigrams, create a co-occurence graph representation and after clustering assess the type of change per word by comparing the results against an intersection table. Their motivation is not only to use a graph-based approach, but to extend the approach by enabling change detection for all parts of speech as opposed to the original model.

\section{Results and Discussion}
Table \ref{table:tab1} shows the results of the shared task. All teams receive better results than baseline 1 (FD), of which a total of 8 teams outperform baseline 2 (CNT + CI + CD). The 4 top scores with $\rho$ $>$ 0.7 are either modified versions of SGNS + OP + CD or use SGNS + VI + CD. The following 4 scores in the range of 0.5 $<$ $\rho$ $<$ 0.6 are generated by the models fastText + OP + CD, SGNS + OP + CD/JSD, and PPMI + WI + CD. 

Contrary to the results by \citet{DBLP:journals/corr/abs-1906-02979} the modified version of vector initialization shows high performance similar to OP alignment, as previously reported by \citet{hamilton-etal-2016-diachronic}. Some modifications to the SGNS + OP + CD approach are able to yield better results than others, e.g. noise-aware alignment and binarized matrices as compared to frequency-driven OP alignment or local alignment with KNN. Team SnakesOnAPlane compare two distance measures and their results show that JSD ($\rho$ $=$ .561) performs minimally worse than CD ($\rho$ $=$ .565) as the semantic change measure for their model.

The overall best-performing model is Skip-Gram with orthogonal alignment and cosine distance (SGNS + OP + CD) with similar hyperparameters as in the model architecture described previously \citep{DBLP:journals/corr/abs-1906-02979}. Said architecture was used as the basis for the two best performing models. Team tidoe reports that binarizing matrices leads to a generally worse performance ($\rho$ $=$ .811) compared to the unmodified version of SGNS + OP + CD ($\rho$ $=$ 0.9). The noise aware alignment approach applied by team sorensbn obtains a higher score ($\rho$ $=$ .854) compared to the result reported by tidoe, but is unable to exceed the performance of the unmodified SNGS + OP + CD for the same set of hyperparameters (window size = 10, negative sampling = 1; subsampling = None). Of the 8 scores above the second baseline, 5 use an architecture that builds upon SGNS + OP + CD. Whereas in the lower score segment $\rho$ $<$ 0.5 none of the models use SGNS + OP + CD. These findings are in line with the results reported by \citet{DBLP:journals/corr/abs-1906-02979}, however the overall best results are lower in this shared task, which is expected from the smaller number of parameter combinations explored.  Additionally, in the shared task the objective was to report the best score and not to calculate the mean which makes it more difficult to compare the robustness of the models presented here.

\bibliography{acl2019}
\bibliographystyle{acl_natbib}

\end{document}